### v1

# Learning Prehensile Dexterity by Imitating and Emulating State-only Observations

Abstract
When humans learn physical skills (e.g., learn to play tennis), we tend to first observe and learn what an expert is doing. But this is often insufficient. Therefore, we subsequently engage in practice, where we try to emulate the expert. Inspired by this observation, we introduce Combining IMitation and Emulation for Motion Refinement (CIMER) – a two-stage framework to learn dexterous prehensile manipulation skills from state-only observations. CIMER’s first stage involves imitation: simultaneously encode the complex interdependent motions of the robot hand and the object in a structured dynamical system. This results in a reactive motion generation policy that provides a reasonable motion prior, but lacks the ability to reason about contact effects due to the lack of action labels. The second stage involves emulation: learn a motion refinement policy to make adjustments to the motion prior of the robot hand such that the desired object motion is reenacted. CIMER is both task-agnostic (no task-specific reward design or shaping) and intervention-free (no need for additional teleoperated or labeled demonstrations). Detailed experiments reveal that i) Imitation alone is insufficient, but adding emulation drastically improves performance, ii) CIMER outperforms existing methods in terms of sample efficiency and the ability to generate realistic and stable motions, iii) CIMER can either zero-shot generalize or learn to adapt to novel objects from the YCB dataset, even outperforming expert policies trained with action labels in most cases.
I Introduction
Learning dexterous manipulation skills involving multi-finger hands presents numerous challenges, such as high-dimensional state and action spaces, complex nonlinear dynamics, and contact effects. Unfortunately, when employed alone, both imitation learning (IL) and reinforcement learning (RL) tend to struggle with such complex high-dimensional tasks. Specifically, IL suffers from distribution shift [1, 2] and RL from poor sample efficiency [3]. To address these challenges, prior works have had success in leveraging demonstrations to expedite reinforcement learning [4, 5, 6, 7, 8, 9].
However, collecting demonstrations for dexterous manipulation in the form of state-action pairs can be challenging and cumbersome due to complex system design and setup [10, 11]. To alleviate this burden, recent research has focused on learning dexterous manipulation skills from state-only observations [12, 13, 14]. However, learning purely from action-free data can result in kinematic motions that are oblivious to force and contact. This lack of knowledge is particularly limiting in dexterous prehensile manipulation tasks (e.g., grasping and tool use) as they exhibit heightened sensitivity to the applied force [12, 15, 16, 17]. To address this limitation, existing methods often require additional teleoperation demonstrations [14, 18], human-in-the-loop corrections [19], task-specific rewards [5, 13, 20, 21], or user-defined sub-goal images [7, 22].
In this work, we contribute a novel task-agnostic framework to learn dexterous prehensile manipulation skills from state-only observations in two stages. We call this framework CIMER, short for Combining IMitation and Emulation for Motion Refinement (see Fig. 1). The first stage is Imitation: CIMER encodes the interdependent desired motions of both the robot hand and the object from the observations into a time-invariant dynamical system that acts as reactive motion generation policy. The second stage is Emulation: CIMER learns a motion refinement policy to adjust the generated robot motion by optimizing a task-agnostic reward that compensates for the lack of action information.
Figure 1: CIMER is a task-agnostic and intervention-free framework to learn dexterous manipulation skills from state-only observations by learning to first generate interdependent desired motions of the robot hand and the object (Imitation), and then refine the generated robot motion in order to reenact the desired object motion (Emulation).
Three key insights motivate our design of CIMER. First, dexterous prehensile manipulation involves complex and inherently interdependent motions of the robot hand and the object. As such, CIMER’s imitation stage captures both their motions simultaneously using a structured dynamical system (see Sec. III-C). Second, state-only observations of successful task executions contain valuable information about the robot hand’s motion, and importantly, unambiguously demonstrate how the object is supposed to move. The lack of action labels only impacts the motion of the robot hand, but not the object. As such, CIMER learns to refine the robot hand motion via reinforcement to ensure that the refined hand motion emulates the desired object motion (see Sec. III-D). Third, separating motion generation and refinement amounts to first learning what one must do (which can be done from state-only observations) and then learning how to do it (which can be learned by self-guided experimentation). Hence, unlike prior methods, CIMER neither requires in-context teleoperated demonstrations (intervention-free) nor task-specific reward designs (task-agnostic). We refer the reader to Sec. II for a detailed and nuanced discussion of related work.
To both experimentally investigate the central issues and to evaluate CIMER, we conducted a series of thorough experiments on three dexterous prehensile manipulation tasks: Tool Use, Object Relocation, and Door Opening.
First, we demonstrate that directly tracking motions learned from state-only observations (i.e., without refinement) using a standard PD controller is rarely successful and necessitates meticulous expert tuning when it does. But CIMER achieves significant improvements in task performance, even while using the same PD controller across different tasks. We also uncover qualitative insights into how CIMER refines motions and why they succeed.
Second, we demonstrate that CIMER tends to achieve better sample efficiency compared to existing methods that learn purely from state-only observations. Moreover, we highlight CIMER’s ability to effectively leverage motion priors from observations to produce realistic behaviors, in contrast to baselines that tend to exploit the simulator and generate unrealistic behaviors.
Third, we show that CIMER significantly outperforms the baselines in terms of zero-shot generalization to unseen objects from the YCB dataset [23]. Notably, CIMER surpasses even the expert policy trained with action labels in many instances. We also demonstrate that CIMER is significantly faster in learning to work with new objects (i.e., skill transfer) for which it initially lacks good zero-shot generalization.
In summary, we contribute a task-agnostic and intervention-free framework called CIMER to learn dexterous prehensile manipulation skills from state-only observations. CIMER is task-agnostic as it can learn to effectively manipulate a variety of objects without relying on bespoke task-specific rewards. CIMER is intervention-free since it does not rely on in-situ expert demonstrations or teleoperation. Further, CIMER learns to ultimately produce PD targets instead of torques, which both promotes safety and facilitates hardware deployment as most multi-fingered hands only support position-based control.
II Related Work
In this section, we discuss our contributions within the context of different related bodies of work.
Learning dexterity from videos: Given the abundance of video data on the internet and the ease of data collection, recent research has focused on learning dexterous manipulation skills from videos. Indeed, videos are a natural source of state-only observations. A key challenge in learning from videos or state-only observations is the lack of data that speaks to the interaction of the robot and the object or the environment. Existing works circumvent this challenge in one of two ways. First, robots can be given an opportunity to learn from interactions, either in simulation [5, 9, 24, 25] or in the real world [7, 8, 22], in addition to videos. Second, robots can leverage additional in-domain teleoperated demonstrations from an expert [6, 14, 18, 20, 26]. Our work falls in the first category. Some of these prior works are limited to replicating a specific and constrained hand motion from a video clip, such as object grasping [8, 9, 24, 25], bagel flipping [6] or bottle opening [26]. While the other methods are capable of generating a variety of desired hand motions and can handle various object targets, they require one of the following: task-specific reward engineering [5, 20], user-provided sub-goal images as reward signals [7, 22], refinements tailored for dexterous grasping [12, 27], additional teleoperated robot demonstrations [14, 18]. In contrast, CIMER learns to generate a wide range of desired hand motions, without relying either on task-specific reward designs or expert intervention.
Dexterity from teleoperation: A challenge related to learning from state-only observations is that of enabling a human operator to teleoperate a dexterous robotic hand [11, 28, 29, 30]. However, it is important to note that these methods are not learning autonomous policies. Instead, they leverage the real-time corrections and guidance provided by the human operator who monitors the robot’s behavior and compensates for its lack of knowledge about contact effects [31].
Learning low-dim skills from videos: There is a large body of work that focuses on learning polices for low-dimensional manipulators (serial-link robots with parallel jaw grippers) from videos. One group of methods utilize a hierarchical framework to learn task-specific waypoint trajectories from human videos that can then be tracked using standard controllers [32, 33, 34, 35]. Another group learn a high-level planner that operates in a latent space and then use it to condition a low-level planner that can produce diverse robot behaviors [36, 37, 38]. However, unlike the low-dimensional skills that are handled by these methods, dexterous manipulation with multi-fingered hands inherently involves more complex finger-object interactions and contact effects. Further, similar to their counterparts that learn dexterous skills, many of these methods also rely on additional teleoperated demonstrations.
Learning locomotion from observations: Our technical approach is inspired by research in learning locomotion skills. Prior work has demonstrated that robots can acquire locomotion skills through imitating real animals [39] or animated characters [40, 41]. These methods first extract the reference motion from video clips, and to account for foot-ground contact effects, refine the learned motions such that the resulting gaits closely track the demonstrated ones. However, given the object-centric nature of dexterous manipulation tasks, a reward strategy centered solely on imitating the observed robot motion may not necessarily ensure task success. This is precisely why CIMER involves an emulation stage where we refine the robot hand motion in order to track the desired object motions.
III Approach
We begin by formulating the problem of learning dexterous manipulation skills from state-only observations, and then supply the details of our proposed solution.
III-A Problem Formulation
Let $\mathcal{D}=[\{{\mathrm{h}}_{t}^{(1)},{\mathrm{o}}_{t}^{(1)}\}^{t=T^{(1)}}_{t=1},\cdots,\{{\mathrm{h}}_{t}^{(N)},{\mathrm{o}}_{t}^{(N)}\}^{t=T^{(N)}}_{t=1}]$ denotes a dataset of $N$ state-only observations of a prehensile manipulation skill, where ${\mathrm{h}}_{t}^{(n)}\in\mathcal{H}\subseteq\mathbb{R}^{n}$ and ${\mathrm{o}}_{t}^{(n)}\in\mathcal{O}\subseteq\mathbb{R}^{m}$ respectively denote the robot hand state and object state (e.g., hand joint positions and object 6D poses) at time $t$ of $n$-th observation. We focus on prehensile manipulation skills since they are particularly challenging to learn without action labels. (see Sec. IV-B). Our goal is to learn a policy from the dataset $\mathcal{D}$ that will enable the robot to autonomously perform the associated dexterous skill. In addition to the dataset, we assume access to a simulator. But, to ensure a task-agnostic solution, we do not allow task-specific reward design or shaping. To remain intervention-free, we do not assume access to additional labeled demonstrations. As such, our problem requires an approach that learns a robust dexterous manipulation skill purely from state-only observations and self-guided interactions with a simulator.
III-B Solution Overview
Our framework to address the above problem, Combining IMitation and Emulation for Motion Refinement (CIMER), operates in two stages: i) Imitation: learn a Motion Generation Policy $\Phi$ in the form of a structured dynamical system that encodes the desired reference motions for both the robot and the object from state-only observations, and ii) Emulation: learn a Motion Refinement policy $\Psi$ that compensates for the lack of action information by adjusting the generated reference motions such that the desired object motion is reenacted. The overall pseudo-code is given in Appendix  C.
III-C Motion Generation via Imitation
We view the complex interdependent desired motions of the robot and the object as solutions to a learnable underlying nonlinear behavioral dynamical system [2, 42, 43]. In recent work, we developed a framework named KODex [44] that learns such underlying dynamics in a highly computationally-efficient manner using Koopman operator theory. Building on this, CIMER employs Koopman operator theory to learn the Motion Generation Policy $\Phi$ from the dataset $\mathcal{D}$. While KODex relies on demonstrations with action labels, CIMER does not. Formally, we formulate the $\Phi$ as a linear dynamical system in the lifted space: $g({\mathrm{h}}_{t+1},{\mathrm{o}}_{t+1})=\mathbf{K}\ \phi({\mathrm{h}}_{t},{\mathrm{o}}_{t})$, where $\phi:\mathbb{R}^{n+m}\rightarrow\mathbb{R}^{p}$ (with $p>(n+m)$) is a user-defined function that “lifts” the robot hand and object states to a higher-dimensional space in which the underlying dynamics appear linear [45], and $\mathbf{K}$ is the Koopman matrix. We use a task-agnostic second-order polynomial lifting function for all tasks in all our experiments. A key benefit of this approach is that we can use least squares to analytically compute $\mathbf{K}$ from $\mathcal{D}$ such that $\Phi$ effectively predicts the time evolution of the lifted states. To extract the original states without a decoder, we augment the original states to the lifted states before learning the Koopman operator. See Appendix D for details. We then use the learned dynamical system as an autonomous policy to generate reference hand motions for any initial condition.
III-D Motion Refinement via Emulation
Merely tracking the hand trajectory generated by the motion generation policy $\Phi$ does not guarantee task success in dexterous prehensile manipulation tasks due to the small margin of error. Even slight deviations from the expert hand motion can have a drastic impact on the object (see Sec. IV-B for empirical validation). We address this issue by letting the robot “practice” the skill learned during imitation by emulating the desired object motion. Specifically, CIMER learns to refine finger and palm trajectories such that the robot reenacts the desired object motion. Interacting with the object helps compensate for the lack of action labels and account for contact effects. Once learned, CIMER employs a task-agnostic PD controller to track the refined hand motions.
We formulate emulation as a reinforcement learning problem. We train a Motion Refinement Policy $\Psi$ to maximize a reward function that captures the essence of emulation: reenacting the desired object motion. To reduce the variance in the training, we use Generalized Advantage Estimator (GAE) [46] and PPO [47].
Figure 2: CIMER refines the motion prior based on the context.
State & Action Space: We parameterize $\Psi$ as an MLP network that predicts the current action: $\hat{\mathrm{h}}_{t}=\Psi(\mathrm{q}_{t},\mathrm{g}_{t})$. Here, $\hat{\mathrm{h}}(t)\in\mathcal{H}$ contains the refined PD targets of the robot hand at time $t$, $\mathrm{q}_{t}=(\bar{\mathrm{h}}_{t-2:t},\bar{\mathrm{o}}_{t-2:t},\hat{\mathrm{h}}_{t-3:t-1})$ is the context containing the history of hand and object states and refined PD targets, and $\mathrm{g}_{t}=(\mathrm{h}_{t+1,t+5,t+10},\mathrm{o}_{t+1,t+5,t+10})$ provides the motion prior (predicted by the motion generation policy $\Phi$) that needs to be refined. This design allows CIMER to reason about action effects when refining the hand motion generated by the motion generator. See Fig. 2 for an illustration.
Rewards: Unlike prior works that rely on task-specific reward design and shaping, we use tracking errors of both the robot hand and the object as reward signals. Formally, we define the reward function as $r_{t}=r_{t}^{h}+r_{t}^{o}+r_{t}^{b}$, where ${r_{t}^{h}=\exp\left(-k^{h}||\bar{\mathrm{h}}_{t+1}-\mathrm{h}_{t+1}||^{2}\right)}$, ${r_{t}^{o}=\exp\left(-k^{o}||\bar{\mathrm{o}}_{t+1}-\mathrm{o}_{t+1}||^{2}\right)}$, $r_{t}^{b}$ is a bonus reward that is trigger when the object tracking error is smaller than a threshold $\epsilon_{o}$. Here, $\bar{\mathrm{h}}_{t+1},\mathrm{h}_{t+1}$, and $\bar{\mathrm{o}}_{t+1},\mathrm{o}_{t+1}$ denote the hand and object states from both the robot and reference motion, respectively. Note that $r_{t}^{h}$ minimizes deviation from the reference hand motion $\mathrm{h}_{t+1}$ while $r_{t}^{o}$ encourages the robot hand to emulate the reference object motion $\mathrm{o}_{t+1}$.
Warm Start: Since the motion generation policy $\Phi$ provides a reasonable motion prior, we expect the motion refinement policy $\Psi$ to only make local adjustments and account for contact effects. As such, we use the reference motions from $\Phi$ to warm start $\Psi$ by minimizing a behavior cloning loss. To counter covariate shift, we inject random noise into the policy’s input and use the same outputs as corrective labels [48, 49, 50]. This warm start represents yet another way in which the motion prior influences the emulation process (in addition to via the reward function and the policy input).
IV Experimental Evaluation
In this section, we report results that i) validate the need for combining imitation with emulation (see Sec. IV-B), ii) demonstrate that CIMER exhibits better sample efficiency compared to baselines and that it can generate more realistic behaviors (see Sec. IV-C), and iii) show that CIMER significantly outperforms baselines, and even the expert in some cases, in terms of both zero-shot generalization and skill transfer to novel objects (see Sec. IV-D). Please refer to the appendices for various additional details.
IV-A Experimental Design
Evaluation Platform: We conducted all our experiments on the widely-used ADROIT Hand [4] – a 30-DoF simulated system (24-DoF hand + 6-DoF floating base) built with MuJoCo Simulator [51].
Data and Expert Policy: We utilized expert DAPG policies [4] to generate 200 state-only observations for each task, comprising both hand and object states. Note that these expert policies were trained using action labels.
Tasks: We consider three prehensile manipulation tasks originally proposed in [4] (see Fig. 3 and Appendix A).
Figure 3: We evaluate on three dexterous prehensile skills from [4]: Tool Use (left), Object Relocation (center), and Door Opening (right).
  • Tool Use: Pick up the hammer to drive the nail into the board placed at a randomized height.
  • Object Relocation: Move an object to a randomized target location (green sphere).
  • Door Opening: Given a randomized door position, unlock the latch and pull the door open.
Metrics: We quantify performance in terms of Task success rate (see Appendix B for criteria) and sample efficiency, always reported across five random seeds unless specified.
IV-B Why Care about Emulation?
We first demonstrate why it is not sufficient to merely imitate observations without emulation. Specifically, the drawbacks of merely using a PD controller to track the hand motions generated by the motion generation policy $\Phi$. To this end, we compare CIMER against two baselines: a). Expert Obs. + PD : Observed expert hand motion tracked using a PD controller, and b). Motion Gen. + PD: Reference hand motion (generated by $\Phi$) tracked using a PD controller. We employed the same PD tracking controller with identical PD gains for all three methods and three tasks, and report the task success rate over the 200 initial conditions for each task.
TABLE I: Tracking state trajectories from an expert or an imitation policy is insufficient for prehensile manipulation. Combing imitation and emulation (CIMER) dramatically improves task performance.
Policy Task Task Success Tool Relocation Door
Rate
Expert Obs. + PD 28.0% 34.5% 10.5%
Motion Gen. + PD 32.0% 35.5% 10.5%
CIMER + PD 92.7($\pm$1.7)96.0($\pm$0.5)76.3($\pm$13.4)%
Figure 4: Object tracking is significantly more important than hand tracking for refining hand motions during emulation.
Figure 5: Intuitive refinements emerge from CIMER’s emulation. Top: Hand ensures hammer hits closer to nail’s center, and applies larger driving force; Bottom Left: Fingers exert more force to ensure firmer grasps and stable transport; Bottom Right: Hand rotates faster to boost momentum while turning door handle (enclosed by dotted lines).
Figure 6: Top: CIMER is considerably more sample efficient than the baselines in two tasks (Tool Use and Object Relocation tasks). The door opening task is an exception likely due to larger refinements being necessary to firmly grasp and swing the door open. Bottom: Qualitative analyses reveals that both the baselines exploit the simulator and generate unrealistic and aggressive behaviors in all three tasks (see dashed red circles indicating loss of grasp, bouncing objects, and hand-object penetration). In contrast, CIMER generates realistic motions as its refinement is anchored by the motion priors from its imitation phase.
As seen in Table I, CIMER significantly outperforms both Expert Obs. + PD nor Motion Gen. + PD. In fact, CIMER is three to four times more successful. These results support our claim that merely tracking an observed trajectory or imitating expert observations are insufficient to perform prehensile manipulation tasks. This is likely due to the fact that contact effects and grasp forces are ignored and inconsequential when merely trying to replicate an expert’s hand motions. In contrast, introducing emulation allows the robot to “practice” what is learned during imitation, making necessary refinements to reenact the desired object motion.
Another noteworthy observation is that to obtain such significant task improvement in the Door Opening task indeed requires another robot DoF for the motion refinement. Our conjecture is that refining only the robot finger and palm motions (24 DoFs) may not suffice for this particular task. Consequently, for this task, we augment the motion refinement policy by including another DoF of the 6-DoF floating base, which is the reference rotational motion along the hand shaft axis. This adjustment potentially enables the hand to generate larger momentum for effectively turning the handle, thereby leading to further task improvement. The task success rate for this task is 46.5($\pm$3.4)%, if only refining the robot finger and palm motions.
A key aspect of our approach to emulation is the increased focus on the object. To validate this focus, we evaluated the relative importance of hand- and object-tracking rewards. Results in Fig. 4 highlight the crucial role of object tracking rewards. Importantly, using hand tracking rewards alone results in little to no performance gains.
To better understand how emulation improves performance, we qualitatively analyzed the modifications made by the motion refinement policy to compensate for the lack of action labels. As shown in Fig. 5, for the Tool Use task, employing the motion refinement policy results in the hammer hitting the nail closer to its center point, and the force of impact (measured by a touch sensor on the nail head) is also increased. As a result, the nail can be fully driven into the board, leading to improved task performance. For the Object Relocation task, the grasping force applied from each fingertip on the object is noticeably increased, indicating a secure grasp during the initial contact and transportation towards the target position. For the Door Opening task, we observe that the hand initially rotates in one direction, then swiftly changes direction to generate greater momentum for turning the handle (the dashed lines denote the handle turning phase). Therefore, we find that emulation indeed nudges the robot to apply the necessary forces and adjusts the hands’ motion so that the object is manipulated as desired. Finally, it’s important to note that these analyses are only possible due to the hierarchical structure of our framework, in contrast to other end-to-end methods.
IV-C Comparisons to Baselines
No existing framework can directly accommodate all the constraints and challenges of our problem formulation (see Sec. III-A). However, we compare CIMER to the following baselines to demonstrate its relative benefits:
  • Pure RL: We adopt TRPO [52] as a pure RL baseline. This baseline is trained without any offline data, but with per-task reward engineering [4].
  • SOIL: A SOTA state-only imitation method for dexterous manipulation [13] that learns an inverse dynamics model to predict actions from state-only observations. SOIL also requires per-task reward engineering [4].
For SOIL and Pure RL methods, we utilized the same model architectures and training parameters as recommended in their original works, and we include the training parameters for CIMER in Appendix F. For quantify performance in terms of task success rates, with means and standard deviations computed across five random seeds.
As shown in Fig. 6 (top), CIMER exhibits significantly better sample efficiency than SOIL and Pure RL for both Tool Use and Object Relocation tasks. In particular, while the baselines struggle to make any progress on the Object Relocation task, CIMER achieved 100% task success rate. In contrast, baselines were able to achieve higher success rates in the Door Opening task with fewer samples compared to CIMER. This is be partly due to the door opening task likely requiring CIMER to make more significant motion refinements to generate enough momentum that will swing the door open.
A closer examination reveals something important: the baselines tend to exploit the simulator and generate unrealistic or unsafe motions across all tasks. The exemplar rollouts in Fig. 6 (bottom) show that the baselines drive the hand to “throw” the hammer, hit the ball, and penetrate the door handle, while CIMER generates more realistic and stable motions. This is because CIMER’s motion generation policy employs a structured dynamical system that provides a natural and effective motion prior for the motion refinement policy to adjust. These results also suggest that CIMER is better suited for real-world deployment and will not require significant efforts in ensuring safety and efficacy.
IV-D Generalization to new objects
In addition to the objects considered thus far, we evaluated CIMER and the baselines on their ability to generalize to 17 novel objects (3 artifacts and other 14 from the YCB dataset [23]) in the Object Relocation task (see Fig. 7). Note that the Ball is the only default object used for training the expert policy with action labels and for collecting state-only demonstrations.
IV-D1 Zero-shot Generalization
Figure 7: We trained candidate policies on one object, and evaluated their ability to generalize and adapt to 17 novel objects.
We report the zero-shot generalization performance of each method and the expert across five random seeds in Fig. 8. Given that both the baselines struggled to learn to relocate the default Ball object, it is no surprise that they consistently fail to generalize to novel objects. While the expert achieves a 100% success rate on the default object (ball) with the benefit for action labels, CIMER notably outperforms the expert on the majority of the novel objects. This impressive ability to readily generalize to novel objects could be attributed to CIMER’s hierarchical structure and suggests that CIMER’s motion prior and refinement strategies are robust to changes in the object’s geometric and kinematic properties. Furthermore, CIMER demonstrates the most consistent performance across different training seeds, as indicated by shorter error bars.
Figure 8: CIMER provides impressive zero-shot generalization to novel objects, even outperforming the expert policy in most cases
IV-D2 Skill Transfer to New Objects
We also evaluated if transferring the skills learned on the default object (Ball) to novel objects would improve sample efficiency (see Appendix G). We fine-tuned CIMER’s motion refinement policy on six novel objects, each of which resulted in 50% or lower success rate during zero-shot generalization. We compared against a similarly-finetuned expert policy [4], and two policies trained from scratch (one using CIMER and the other using the expert method).
Figure 9: Skill transfer provides a significant boost to CIMER’s sample efficiency (top two rows), unless the target object is considerably different from the source (bottom row). Solid lines show mean trends and shaded areas show $\pm$ standard deviation over five random seeds.
As shown in Fig. 9, transferring motion refinement skills considerably improves task performance in four of the six objects. On these four objects, CIMER with skill transfer outperforms the baselines trained from scratch as well as the fine-tuned expert. In contrast, skill transfer seems to hurt performance when dealing with other two objects. This could be because skill transfer is counter-productive when target objects are drastically different from the source object, requiring a very different grasping strategy. Once again, CIMER outperforms the expert method on all these novel objects, if trained from scratch or fine-tuned.
V Limitations and Future Work
Our work represents the first task-agnostic and intervention-free approach to learning dexterous prehensile manipulation skills from state-only observations. But it leaves behind several avenues for further improvement. First, given that human videos are ubiquitous online, one could integrate recent advances in motion retargetting [5] with CIMER to directly learning from videos. Second, we do not consider the manipulation of deformable and fragile objects. It is yet unclear how to represent such objects and incentivize emulation for such objects. Third, some objects require a completely different hand trajectory (e.g., grasping a pan by the handle vs. a ball). It might be possible to enable such drastic adaptations by refining the motion of the hand base in addition to the fingers and wrist. Fourth, estimating local contacts using tactile sensors [53, 54, 55] might help refine the hand motion. Last, CIMER is yet to be implemented on physical robots. Leveraging the fact that CIMER outputs PD targets, we plan to deploy on hardware and validate its benefits in the real-world.
Appendix A Task State Space Designs
The state design for each task is as follows:
Tool use: For this task, the floating wrist base can only rotate along the $x$ and $y$ axis, so we have $\mathrm{h}_{t}\in\mathcal{H}\subseteq\mathbb{R}^{26}$. Regarding the object states, unlike the other tasks, where the objects of interest are directly manipulated by the hand, this task requires to modify the environment itself. As a result, except for the hammer positions, orientations and their corresponding velocities $\mathrm{p}^{\text{tool}}_{t},\mathrm{o}^{\text{tool}}_{t},\dot{\mathrm{p}}^{\text{tool}}_{t},\dot{\mathrm{o}}^{\text{tool}}_{t}$ ($\mathbb{R}^{3}$), we also define the nail goal position $\mathrm{p}^{\text{nail}}$ ($\mathbb{R}^{3}$). Finally, we have $\mathrm{o}_{t}=[\mathrm{p}^{\text{tool}}_{t},\mathrm{o}^{\text{tool}}_{t},\dot{\mathrm{p}}^{\text{tool}}_{t},\dot{\mathrm{o}}^{\text{tool}}_{t},\mathrm{p}^{\text{nail}}]\in\mathcal{O}\subseteq\mathbb{R}^{15}$. We use the $\mathrm{p}^{\text{tool}}_{t}$ and $\mathrm{o}^{\text{tool}}_{t}$ as the object tracking states.
Object relocation: For this task, the ADROIT hand is fully actuated, so we have $\mathrm{h}_{t}\in\mathcal{H}\subseteq\mathbb{R}^{30}$ (24-DoF hand + 6-DoF floating wrist base). Regarding the object states, we define $\mathrm{p}^{\text{target}}$ and $\mathrm{p}^{\text{ball}}_{t}$ as the target and current positions. Then, we compute $\bar{\mathrm{p}}^{\text{ball}}_{t}=\mathrm{p}^{\text{ball}}_{t}-\mathrm{p}^{\text{target}}$, which is the component of $\mathrm{p}^{\text{ball}}_{t}$ in a new coordinate frame that is constructed by $\mathrm{p}^{\text{target}}$ being the origin. We additional include the ball orientation $\mathrm{o}^{\text{ball}}_{t}$ and their corresponding velocities $\dot{\mathrm{p}}^{\text{ball}}_{t}$, $\dot{\mathrm{o}}^{\text{ball}}_{t}$ (all $\mathbb{R}^{3}$). Finally, we have $\mathrm{o}_{t}=[\bar{\mathrm{p}}^{\text{ball}}_{t},\mathrm{o}^{\text{ball}}_{t},\dot{\mathrm{p}}^{\text{ball}}_{t},\dot{\mathrm{o}}^{\text{ball}}_{t}]\in\mathcal{O}\subseteq\mathbb{R}^{12}$. We use the $\bar{\mathrm{p}}^{\text{ball}}_{t}$ as the object tracking states.
Door opening: For this task, the floating wrist base can only move along the direction that is perpendicular to the door plane but rotate freely, so we have $\mathrm{h}_{t}\in\mathcal{H}\subseteq\mathbb{R}^{28}$. Regarding the object states, we define the fixed door position $\mathrm{p}^{\text{door}}$, which can provide with case-specific information (similar to $\mathrm{p}^{\text{nail}}$ in Tool Use task), and the handle positions $\mathrm{p}^{\text{handle}}_{t}$ (both $\mathbb{R}^{3}$). In order to take into consideration the status of door being opened, we include the angular velocity of the opening angle $v_{t}$$(\mathbb{R}^{1})$. Finally, we have $\mathrm{o}_{t}=[\mathrm{p}^{\text{handle}}_{t},v_{t},\mathrm{p}^{\text{door}}]\in\mathcal{O}\subseteq\mathbb{R}^{7}$. We use the $\mathrm{p}^{\text{handle}}_{t}$ as the object tracking states.
For all tasks, the simulation ran at 500HZ. Note that for CIMER and baselines, the robot states remain the same, whereas for object states, we adhere to the design choices proposed in the original works for the baselines [4, 13].
Appendix B Task Success Criteria
We used the same task success criteria as in [4, 13, 44]:
Tool Use: The task is considered successful if at last time step $T$, the Euclidean distance between the final nail position and the goal nail position is smaller than 0.01.
Object Relocation: At each time step $t$, if $\sqrt{|\mathrm{p}^{\text{target}}-\mathrm{p}^{\text{ball}}_{t}|^{2}}<0.10$, then we have $\rho(t)=1$. The task is considered successful if $\sum_{t=1}^{T}\rho(t)>10$.
Door Opening: The task is considered successful if at last time step $T$, the door opening angle is larger than 1.35 rad.
Learning Prehensile Dexterity by Imitating and Emulating State-only Observations
Yunhai Han, Zhenyang Chen, Harish Ravichandar
Georgia Institute of Technology
Email: {yhan389,zchen927,harish.ravichandar}@gatech.edu
Appendix C CIMER Pseudo-code
The overall pseudo-code for CIMER is given below.
Inputs: State-only observation dataset $\mathcal{D}$, Koopman lifting function $g(\cdot)$;
Motion Imitation
1 Formulate Motion Generation Policy $\Phi$ as a Koopman-based dynamical system: $g({\mathrm{h}}_{t+1},{\mathrm{o}}_{t+1})=\mathbf{K}\ g({\mathrm{h}}_{t},{\mathrm{o}}_{t})$
2 Regress Koopman Matrix $\mathbf{K}$ on $\mathcal{D}$;
Motion Emulation
3 Warm-start the Motion Refinement Policy $\Psi$ to reconstruct trajectories generated by $\Phi$;
4 for $iter\in\{1,...,\max\}$ do
5       Initialize replay buffer $\mathcal{B}=\varnothing$ ;
6       for $k\in\{1,...,M\}$ do
7             Specify random initial states $\{{\mathrm{h}}_{1}^{(k)},{\mathrm{o}}_{1}^{(k)}\}$;
8             Generate reference motions $\{{\mathrm{h}}_{t}^{(k)},{\mathrm{o}}_{t}^{(k)}\}^{t=T^{(k)}}_{t=2}$ by rolling out $\Phi$;
9             for $t\in\{1,...,T^{(k)}-1\}$ do
10                   Motion Refinement Policy $\Psi$ inputs $\mathrm{s}_{t}$, and outputs $\mathrm{a}_{t}$ for execution;
11                   Compute the tracking reward $r_{t}$;
12             end for
13            Add $\{\mathrm{s}_{t}^{(k)},\mathrm{a}_{t}^{(k)},r_{t}^{(k)}\}^{t=T^{(k)}-1}_{t=1}$ to $\mathcal{B}$;
14
15       end for
16      Update the Motion Refinement Policy $\Psi$ with $\mathcal{B}$;
17
18 end for
Return: Trained policies $\Phi,\Psi$
Algorithm 1 CIMER
Appendix D Details of KODex Training
D-A Modeling Dexterous Manipulation Skills
A central principle behind KODex is that the desired behavior of a robot can be represented using a dynamical system. To this end, we define the state at time $t$ as $\mathrm{x}(t)=[{\mathrm{h}_{t}}^{\top},{\mathrm{o}_{t}}^{\top}]^{\top}$, where ${\mathrm{h}_{t}}\in\mathcal{H}\subseteq\mathbb{R}^{n}$ and ${\mathrm{o}_{t}}\in\mathcal{O}\subseteq\mathbb{R}^{m}$ represent the state of the robot and the object, respectively, at time $t$. As such, the dynamical system we wish to capture is
$\mathrm{x}(t+1)=F^{*}(\mathrm{x}(t))$ (1)
where $F^{*}(\cdot):\mathcal{H}\times\mathcal{O}\rightarrow\mathcal{H}\times\mathcal{O}$ denotes the dynamics that govern the interdependent motions of the robot and the object.
A key challenge in learning the dynamical system in (1) is that it can be arbitrarily complex and highly nonlinear, depending on the particular skill of interest. KODex leverages Koopman operator theory to learn a linear dynamical system that can effectively approximate such complex nonlinear dynamics. To this end, we first define the Koopman lifting function $g(\cdot)$ as follows
$g(\mathrm{x}(t))=[{\mathrm{h}_{t}}^{\top},\psi_{h}({\mathrm{h}_{t}}),{\mathrm{o}_{t}}^{\top},\psi_{o}({\mathrm{o}_{t}})]^{\top},\forall t$ (2)
where $\psi_{h}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{\bar{n}}$ and $\psi_{o}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{\bar{m}}$ are vector-valued lifting functions that transform the robot and object state respectively.
Now, we introduce the Koopman matrix $\mathbf{K}$, which approximates the system evolution in the lifted state space as follows
$g(\mathrm{x}(t+1))=\mathbf{K}g(\mathrm{x}(t))$ (3)
D-B Learning the Koopman Matrix $\mathbf{K}$
The next step is to learn the Koopman matrix $\mathbf{K}$ from dataset $\mathcal{D}=[\{{\mathrm{h}}_{t}^{(1)},{\mathrm{o}}_{t}^{(1)}\}^{t=T^{(1)}}_{t=1},\cdots,\{{\mathrm{h}}_{t}^{(N)},{\mathrm{o}}_{t}^{(N)}\}^{t=T^{(N)}}_{t=1}]$ (Section. III-B). Recall that CIMER’s Motion Generation Policy $\Phi$ is formulated as the linear dynamical system shown in (3).
As described in [44], we can efficiently compute the Koopman matrix as $\mathbf{K}=\mathbf{A}\mathbf{G}^{\dagger}$, where $\mathbf{A}$ and $\mathbf{G}$ are shown as follows
$\mathbf{A}=\sum_{n=1}^{n=N}\sum_{t=1}^{t=T^{(n)}-1}\frac{g(\mathrm{x}^{(n)}(t+1))\otimes g(\mathrm{x}^{(n)}(t))}{N(T^{(n)}-1)},\ \mathbf{G}=\sum_{n=1}^{n=N}\sum_{t=1}^{t=T^{(n)}-1}\frac{g(\mathrm{x}^{(n)}(t))\otimes g(\mathrm{x}^{(n)}(t))}{N(T^{(n)}-1)}$ (4)
where $\mathbf{G}^{\dagger}$ denotes the Moore–Penrose inverse [^1] It could be efficiently computed using thescipy.linalg.pinv($\mathbf{G}$) function from Scipy library. of $\mathbf{G}$, and $\otimes$ denotes the outer product.
We use the computed Koopman matrix $\mathbf{K}$ to generate rollouts in the lifted state space. However, we need to obtain the rollouts in the original state states to retrieve the predicted robot and object states. Since we designed $g(\mathrm{x}(t))$ such that robot state $\mathrm{h}_{t}$ and the object state $\mathrm{o}_{t}$ are parts of lifted states in (2), we can easily retrieve the reference motions of both the robot and the object $\{\mathrm{h}_{t},\mathrm{o}_{t}\}$ by selecting the corresponding elements in $g(\mathrm{x}(t))$.
Appendix E Emulation might not be necessary for Non-prehensile Manipulation
We present the task success rate on another non-prehensile manipulation task, In-hand Reorientation, which are as follows: Expert Obs. + PD: 91.0%, Motion Gen. + PD: 69.5%, and CIMER: 72.7($\pm$1.0)%. Contrary to the results presented in Table. I, where Expert Obs. + PD shows poor performance on three dexterous prehensile manipulation tasks, in this scenario, it could achieves satisfactory results. This disparity arises because the challenge of the In-hand Reorientation task primarily pertains to the generation of precise and diverse hand motions, rather than dealing with heightened sensitivity to applied forces. Therefore tracking the observed expert motion directly translates to achieving the desired object motion. This is also why both the Motion Gen. + PD and CIMER are not comparable to Expert Obs. + PD in this In-hand Reorientation task. Therefore, we exclusively apply CIMER policies to the three dexterous prehensile manipulation tasks, aligning with the intended purpose of motion refinement policies. We leave the improvement of motion generation policy for non-prehensile tasks as a topic for future work.
Appendix F CIMER Hyperparameters
Table II summarizes the hyper-parameter settings for training CIMER policies. For the three tasks, the hyper-parameter settings are the same.
TABLE II: Training details
[…]
$(256,128)$ 2e-6 0.98
[…]
0.2 8 0.97
In addition, for all tasks, we used $k^{h}=5,k^{o}=5,r_{t}^{b}=25$ to compute the tracking rewards described in Sec. III-D.
Appendix G Adaptation to New Objects
We also evaluated CIMER’s ability to adapt to novel objects. Specifically, we finetuned the CIMER on each of the seven objects for which its success rates was lower than 50%. To ensure that CIMER remains intervention-free, we did not collect or rely on additional observations. Instead, we reused the old motion generation policy (trained on the default object), and simply changed the object used in the simulator when training the motion refinement policy. We similarly trained the expert policies to ensure a fair comparison.
As shown in Fig. 10, we find that both methods gradually learn to manipulate the new objects, with CIMER exhibiting significantly better sample efficiency almost all objects. The power drill presents a notable exception to this trend, with CIMER particularly struggling to adapt. This is due to the power drill likely requiring a significantly different grasping strategy than the default object (Ball) on which CIMER’s motion generation policy was trained. Indeed, CIMER’s motion refinement only adjusts finger and palm motions, and does not account for grasp synthesis. This could potentially be addressed by also learning to refine the hand base motions with the help of appropriate regularizers [56] that encourage refined motions to remain close to the reference motion.
Figure 10: We compare the RL sample efficiency between CIMER and expert policies when being trained on new objects. For each, solid lines indicate mean trends and shaded areas show $\pm$ standard deviation over five random seeds. We find that both methods will gradually learn the manipulation strategy of the new objects, except for the power drill. Moreover, CIMER policies exhibit the better sample efficiency.